# Semantic Structure-Mapping in LLM and Human Analogical Reasoning

### v1

@@ 3.1.3 Other Observations
3.1.4 Diagnosing the Use of an RHS-Only Heuristic
To clarify whether subjects actually make use of left-right relations or only complete right-side patterns in the Semantic Structure experiment, we design the Relational condition, a $2\times n$ variant of the Defaults condition which cannot be solved (consistently) using only the right-hand terms (see the example in Table 3).
pants => H # H
glove => X # X
torso => V
foot => Z
head => M
shirt => V # V
hat =>
Table 3: An example from the Relational variant of the Defaults task, used to diagnose subjects’ tendency to rely on RHS-only heuristics to solve the task.
Results are shown in Figure 5. Human subjects and Claude 3 exhibit similar performance, with accuracies of approximately 0.7. GPT-4, however, attains much lower accuracy of approximately 0.35. Simple effects analysis shows that GPT-4 obtains significantly worse accuracy than human subjects (coef = -1.3669, z = -3.065, p = 0.002), while the accuracy of Claude 3 does not differ significantly from human subjects (coef = 0.2111, z = 0.467, p = 0.640).
Figure 5: Human and LLM accuracy in the Relational condition followup, with Defaults condition performance for reference. Error bars show standard errors.
@@ 3.1.5 Takeaways
Despite weak performance from many models on our analogical reasoning tasks, GPT-4 and Claude 3 perform well, showing similar patterns to humans in leveraging semantic structure of corresponding domains to solve analogies. However, differences do remain in how they handle semantic structure in the source domain. Humans prefer leveraging semantic structure when a clear pattern exists (evidenced by the Defaults and Random Finals conditions) but can ignore words when structure is lacking (Randoms condition). Models show the former bias but not the latter ability, appearing distracted by random lexical items. Nevertheless, model results increasingly resemble human subjects, suggesting larger models may close this gap.
@@
Furthermore, qualitative differences exist even between the best models. GPT-4 and Claude 3 match human performance in the Defaults condition, but when the structure is generalized from $2\times 2$ to $2\times n$ in the Relational followup, making a right-hand-only strategy unworkable, Claude 3 maintains human-level performance while GPT-4 drops significantly. Despite limited public information, it’s notable that models produced using presumably similar approaches can exhibit meaningfully different behavioral patterns.
3.2 Mapping Semantic Content
@@
The Semantic Structure experiment, which presented subjects with source and target domains with corresponding semantic structure (i.e., with corresponding relations between terms), provides insight into the relative bias of human subjects and models to transfer this structure across domains. The Semantic Content experiment modifies the tasks to investigate the extent to which human subjects and models can transfer elements of the linguistic meaning of terms from one domain to another.
To achieve this, we ensure that elements of the target domain directly depend on properties of corresponding source domain elements, requiring knowledge of the source domain terms’ meaning for perfect performance. As in the Semantic Structure experiment, source and target domains are paired such that patterns in the target domain mirror those in the source domain. Together, these experiments compare the subject’s ability and tendency to use a structure-mapping approach. Four tasks are generated, encoding either one or two dimensions of variation and either involving or not involving numeric reasoning (see Table 4).
Categorial: Right-hand terms are single characters corresponding to a Categorial property of the left-hand terms. chicken => ! spider => ! cat => * horse => * ant => ! dog => * bee => ! human =>
Multi-Attribute: Right-hand terms are a sequence of several characters that vary according to two properties of the left-hand terms. grandfather => ! grandmother => * mother => * * father => ! ! brother => ! ! ! sister =>
Numeric: Right-hand terms are a sequence of a single repeated character, with the number of repetitions corresponding to a numeric property of the left-hand terms. chicken => * * human => * * dog => * * * * spider => * * * * * * * * cat => * * * * horse => * * * * bee =>
Numeric Multi-Attribute: Right-hand terms are a sequence of a repeated character, with the number of repetitions corresponding to a numeric property of the left-hand terms and the character corresponding to a Categorial property. horse => * * * * cat => * * * * ant => ! ! ! ! ! ! bee => ! ! ! ! ! ! chicken => ! ! spider => ! ! ! ! ! ! ! ! dog => * * * * human =>
Table 4: The conditions of the Semantic Content experiment.
Figure 6: Human and model accuracy by condition in the Semantic Content experiment. Error bars show standard errors.
Results for human subjects, GPT-4, and Claude 3 are shown in Figure 6 (other tested models attain much lower accuracy as before).
@@ 3.2.1 Human Performance Continues to be Robust
Human subjects perform robustly and consistently, as in the previous experiment. Human accuracy ranges from 0.4 to 0.8 across conditions, comparable to the earlier Semantic Structure experiment. As expected, subjects generally describe their strategy as relating properties of the left-hand terms to their representations on the right-hand side.
3.2.2 Claude 3 Matches Human Performance Stably Across Conditions
Claude 3 matches human performance stably across the different conditions of the Semantic Content experiment with its accuracy falling into a comparable range of 0.4 to 0.7. The model exhibits marginally better performance in the Multi-Attribute condition and marginally worse performance in the remaining three. These differences are insignificant across all conditions, which covers the Categorial (coef = -0.8109, z = -1.879, p = 0.060), Multi-Attribute (coef = 0.6206, z = 1.478, p = 0.140), Numeric (coef = -0.1788, z = -0.439, p = 0.661), and Numeric Multi-Attribute (coef = -0.2009, z = -0.484, p = 0.629) conditions. Therefore, Claude 3 performs as well as human subjects across all conditions of this experiment.
3.2.3 GPT-4 Lags Human Subjects on Numeric Reasoning
GPT-4 achieves good results in the Categorial and Multi-Attribute conditions, with mean accuracies of approximately 0.7 in both (compared to 0.7 and 0.4 respectively for human subjects). GPT-4 is not significantly worse than humans in the Categorial condition (coef = -0.3927, z = -0.889, p = 0.374). GPT-4 significantly outperforms human subjects in the Multi-Attribute condition (coef = 1.0624, z = 2.429, p = 0.015). However, its accuracy drops to 0.2-0.3 in the remaining conditions and we find that GPT-4 is significantly worse than humans in both the Numeric (coef = -1.4781, z = -3.321, p = 0.001) and Numeric Multi-Attribute conditions (coef = -0.9694, z = -2.185, p = 0.029).
@@ 3.2.3 GPT-4 Lags Human Subjects on Numeric Reasoning
3.2.4 Human Performance Drops in Compositional Conditions, But Models Remain Constant
When comparing the performance of a subject in a non-compositional (single-attribute) condition to the corresponding compositional (multi-attribute) version, we observe some decrease in performance for human subjects but not for models (note that this surprising result is subject to alternative explanations, addressed in the discussion below). The accuracy of human subjects drops from approximately 0.7 to approximately 0.4 when comparing the Categorial condition to the corresponding compositional version (the Multi-Attribute condition). A simple effects analysis confirms that this decline is significant (coef = -1.2267, z = -3.091, p = 0.002). We see a non-significant decrease in accuracy for human subjects when comparing the Numeric condition to its compositional counterpart, with performance dropping from approximately 0.6 to approximately 0.5 (coef = -0.3795, z = -1.028, p = 0.304).
By contrast, we do not find either model to be significantly worse in compositional conditions than non-compositional ones. In fact, GPT-4 exhibits a slight improvement in the compositional conditions, though this change is statistically insignificant for both the Multi-Attribute condition relative to the Categorial condition (coef = 0.2281, z = 0.477, p = 0.634) and for the Numeric Multi-Attribute condition relative to the Numeric condition (coef = 0.1292, z = 0.254, p = 0.799). For Claude 3 we similarly find the differences to be insignificant for the Multi-Attribute condition relative to the Categorial condition (coef = 0.2049, z = 0.452, p = 0.651) and for the Numeric Multi-Attribute condition relative to the Numeric condition (coef = -0.4013, z = -0.893, p = 0.372).
@@ 3.2.5 Takeaways
The Semantic Content experiment confirms that human subjects perform robustly and flexibly across diverse task variations. Claude 3 matches human performance in all conditions, indicating it shares humans’ tendency to use the source domain’s semantic content when completing target domains. While GPT-4’s poor performance in numeric conditions is notable, it reflects a failure in numeric reasoning rather than a difference in analogical reasoning.
We find evidence of decreased human performance, but not model performance, in compositional conditions, contrasting with some existing research [33]. However, other factors may be at play. Models’ negative compositionality effect may be masked by a positive effect, such as increased available information: when the target domain represents two source domain properties, models may more easily recognize the encoding of source domain properties. Human subjects may benefit less from this competing effect if they do not struggle to observe this information encoding.
@@ 4 Discussion
Our results show that the best-performing LLMs are able to successfully complete many analogical reasoning tasks with human-level accuracy using novel stimuli not present in their training data. They also show that there remain meaningful differences in how such analogies are processed, evidenced by differences in how humans and models respond to distracting or misleading information. However, we observe a clear trend: more recent models come increasingly close to matching human performance across our tasks. In particular, Claude 3, the most recently-released model we test, exhibits impressively robust performance across most task variations, even closing the gap with humans in some test conditions in which its predecessor (GPT-4) exhibited limitations (such as the Relational task version in which mapping from the source domain must be used for success). Together, these results raise questions about the ability of LLMs and similar models to serve as candidate cognitive models, which we discuss briefly below.
@@ 4.1 Evaluating the Competence of LLMs
The breadth of Claude 3’s success in our tasks is noteworthy. It suggests that state-of-the-art LLMs can broadly match human performance not only in formal analogical reasoning tasks, as suggested by Webb et al. [4], but also in tasks that require mapping semantic information across linguistic and non-linguistic domains. As such, our results weigh against a long-standing view in cognitive science, according to which connectionist models without a built-in symbolic component are constitutively limited in their ability to robustly handle analogical reasoning tasks [7]. They also inform discussions of whether LLMs possess “functional” linguistic competence, in addition to “formal” linguistic competence [3]. Further work is needed to characterize the precise mechanism that LLMs are using to solve these tasks; it is possible–though increasingly unlikely given the robustness of the behavioral results–that success is due to a myriad of heuristics rather than a systematic analogical reasoning process. Even so, evidence of LLMs completing analogical reasoning tasks in domains designed to involve linguistic structure-mapping, in addition to tasks over abstract symbols, runs counter to the claim that LLMs are capable of formal but not functional linguistic competence.
@@
From this perspective, our results offer evidence to support both sides of the present debate about whether LLMs possess human-level analogical reasoning (see Webb et al. [15], Mitchell et al. [10], and Hodel et al. [16]). Supporting the argument of Webb et al. [15] that deficiencies in capabilities other than analogical reasoning can explain poor model performance in some tasks, we find that GPT-4’s failure in the numeric conditions of our Semantic Content experiment may be due to a deficiency in counting ability. However, contrary to Webb et al. [4], who report impressive analogical reasoning in both GPT-3 and GPT-4, we do find a qualitative difference in the performance of these two models, with GPT-3 performing quite poorly on our tasks. Among the models tested, only GPT-4 and Claude 3 produce results that merit detailed comparison with human subjects. This suggests that claims of human-level performance of LLMs on analogical reasoning tasks may have been premature and might have relied on insufficiently challenging tasks.
@@
Collectively, these patterns bear on the larger question of how we should arbitrate disputes about competence in machine-human comparisons. On the one hand, it seems reasonable to assume that any system that can reliably achieve success at or above human level on experiments like ours–without relying on memorization and other confounds–should be considered competent at analogical reasoning through structure-mapping. On the other hand, we should be open to the possibility that such competence may be implemented differently in LLMs and humans.
@@
4.2 Analogy in Human(-like) Learning and Bootstrapping
@@
Unlike previous research comparing analogical reasoning in human subjects and LLMs, our tasks involve transferring semantic structure and content from source to target domains, rather than reasoning over abstract symbols. Our experiments thus investigate whether LLMs’ analogical reasoning resembles that of human subjects in a manner pertinent to its purportedly central role in broader cognition. Following Gentner [11], emphasis has been placed on relational similarity, rather than just feature similarity, in mapping from a familiar source to a foreign target domain during analogical reasoning to allow for the flexible transfer of knowledge [41, 42, 43]. This conception allows analogical reasoning to play a fundamental role in human cognition, supporting the emergence of diverse cognitive abilities via “bootstrapping” [44, 45, 46]. In bootstrapping, two cognitive processes mutually support each other’s development. In Gentner’s Structure-Mapping Theory (SMT), language development and structure-mapping-based analogical reasoning are hypothesized to co-develop, with structure-mapping developing the necessary relational reasoning to model language-world relations, and language acquisition in turn developing symbolic reasoning capacities that amplify structure-mapping abilities. Consequently, analogical reasoning is seen as a central cognitive phenomenon of interest.
@@
The success of some LLMs in many of our tasks suggests that the most advanced models may be capable of employing a structure-mapping based approach to analogical reasoning, in which relations in the source domain are used to constrain and guide reasoning about relations in the target domain. This raises the possibility that a bootstrapping cycle between language development and analogical reasoning in humans, as proposed by Gentner [44], may be paralleled in language models. The emergence of such competence from training primarily on text prediction would yield new hypotheses about the emergence of analogical reasoning as a central cognitive faculty from generic learning mechanisms (possibly combined with the unique pressures of language acquisition). However, the mixed success of LLMs and the significant differences from humans in certain conditions underscore the need for continued research to test the robustness of any conclusion that analogical reasoning in LLMs closely matches that of human subjects. As LLM outputs continue to converge toward human responses–an expected product of the language modelling objective–it is crucial to develop novel tasks that examine analogical reasoning ability and are not attested in the training data. While our task allows for clear discrimination between human performance and that of most models prior to Claude 3, further differences in analogical reasoning patterns between humans and Claude 3 likely exist beyond those revealed by our tests. More granular testing would help clarify the extent of the remaining discrepancies between humans and the most advanced LLMs, and much further work is required to verify the hypothesis that language models parallel the bootstrapping cycle between language development and analogical reasoning in humans.
The proprietary nature of leading LLMs like Claude 3 unfortunately limits our ability to directly investigate the features that may explain the emergence of a response pattern largely mirroring that of human subjects. However, increasingly sophisticated open-weights models are being released, which may allow for interpretability work to analyze the internal mechanisms of a model and shed light on the underlying mechanisms that enable advanced LLMs to exhibit impressive analogical reasoning abilities in many tasks.
5 Acknowledgments
This work was supported in part by NIH NIGMS COBRE grant #5P20GM10364510.

### v2

@@ 4.1.6 Takeaways
Despite weak performance from many models on our analogical reasoning tasks, GPT-4, Claude 3, and Llama-405B perform well, showing similar patterns to humans in leveraging semantic structure of corresponding domains to solve analogies. However, differences do remain in how they handle semantic structure in the source domain. Humans prefer leveraging semantic structure when a clear pattern exists (evidenced by the Defaults and Random Finals conditions) but can ignore words when structure is lacking (Randoms condition). Models show the former bias but not the latter ability, appearing distracted by random lexical items. Nevertheless, newer models increasingly resemble human subjects, suggesting larger models may close this gap.
@@
Furthermore, qualitative differences exist even between the best models. GPT-4, Claude 3, and Llama-405B match human performance in the Defaults condition, but when the structure is generalized from $2\times 2$ to the $2\times n$ followup, making a right-hand-only strategy unworkable, Claude 3 maintains human-level performance while GPT-4 and Llama-405B drop significantly. Despite limited public information, it’s notable that models produced using presumably similar approaches can exhibit meaningfully different behavioral patterns.
4.2 Reasoning From Semantic Content
@@
The above Semantic Structure experiments provide insight into the relative bias of human subjects and models to transfer this structure across domains. The Semantic Content experiments modify the tasks to investigate the extent to which human subjects and models can transfer elements of the linguistic meaning of terms from one domain to another as elements of the target domain in this experiment directly depend on properties of corresponding source domain elements, requiring knowledge of the source domain terms’ meaning for perfect performance. We again present overall performance observations for humans and LLMs first (§4.2.1), before providing more specific insights about how individual LLMs differ in their performance across conditions (§4.2.2) and about how humans vs. LLMs respond to compositional reasoning tasks (§4.2.3).
Figure 6: Human and model reference match proportion by condition in the Semantic Content experiment. Error bars show standard errors. Outline bars correspond to single-attribute conditions while filled bars correspond to dual-attribute conditions, with color differentiating whether numeric properties or only categorial properties are included. For example, performance in the single-attribute Numeric condition is given by the outlines surrounding peach bars while the peach bars give performance in the Numeric Dual-attribute condition. Note that “Categorial” conditions include one or two categorial properties while “Numeric” conditions include either a numeric property alone or a numeric property paired with a categorial one.
@@ 4.2.1 Human Performance Continues to be Robust
Results for human subjects, GPT-4, and Claude 3 are shown in Figure 6. The other models we tested attain a much lower match with reference answers as before, and thus are omitted from our primary analysis. Human match percentages range from 0.4 to 0.8 across conditions, comparable to the earlier Semantic Structure experiments. In qualitative analysis, subjects generally describe their strategy as relating properties of the left-hand terms to their representations on the right-hand side. For example, one subject describes their strategy as “Looking at other comparable entries to figure out what the answer to the last entry was (dog:puppy::cat:kitten)”, and reporting a strategy in which “I tried to find similar pairs of entries and looked at their meanings.” A.4 contains further qualitative analysis.
4.2.2 LLM-Specific Variations
It’s worth noting that the individual LLMs do not necessarily pattern alike. We describe several interesting idiosyncrasies below.
Claude 3 matches human performance stably across the different conditions of the Semantic Content experiment with its match to reference falling into a comparable range of 0.4 to 0.7. The model exhibits marginally better performance in the Dual-attribute condition and marginally worse performance in the remaining three. These differences are insignificant across all conditions, which covers the Categorial (coef=-0.8109, z=-1.879, p=0.060), Dual-attribute (coef=0.6206, z=1.478, p=0.140), Numeric (coef=-0.1788, z=-0.439, p=0.661), and Numeric Dual-attribute (coef=-0.2009, z=-0.484, p=0.629) conditions. Therefore, Claude 3 performs as well as human subjects across all conditions of this experiment.
GPT-4 achieves good results in the Categorial and Dual-attribute conditions, with mean match proportions of approximately 0.7 in both (compared to 0.7 and 0.4 respectively for human subjects). GPT-4 is not significantly worse than humans in the Categorial condition (coef=-0.3927, z=-0.889, p=0.374). GPT-4 significantly outperforms human subjects in the Dual-attribute condition (coef=1.0624, z=2.429, p=0.015). However, its match to reference answer drops to 0.2-0.3 in the remaining conditions and we find that GPT-4 is significantly worse than humans in both the Numeric (coef=-1.4781, z=-3.321, p=0.001) and Numeric Dual-attribute conditions (coef=-0.9694, z=-2.185, p=0.029).
@@ 4.2.2 LLM-Specific Variations
Llama-405B achieves very consistent results of 0.4-0.5 in three conditions but a lower score of approximately 0.2 in the Numeric Dual-attribute condition. Performance is significantly lower than human subjects in the Categorial condition (coef=-1.3139, z=-3.034, p=0.002), does not differ significantly from human subjects in the Dual-attribute (coef=-0.1904, z=-0.453, p=0.650) or Numeric conditions (coef=-0.6818, z=-1.668, p=0.095), and is significantly worse than human subjects in the Numeric Dual-attribute condition (coef=-1.7346, z=-3.354, p=0.001). Llama-405B therefore struggles to combine reasoning numerically with reasoning about two properties simultaneously in our experiments. This is despite it being able to reason about two non-numeric properties or about a single numeric property.
4.2.3 Effect of compositionality
When comparing the performance of a subject in a non-compositional (single-attribute) condition to the corresponding compositional (Dual-attribute) version, we observe some decrease in performance for human subjects but not for models. This (admittedly unintuitive) result is subject to alternative explanations, which we address in the discussion below. The match to reference answers of human subjects drops from approximately 0.7 to approximately 0.4 when comparing the Categorial condition to the corresponding compositional version (the Dual-attribute condition). A simple effects analysis confirms that this decline is significant (coef=-1.2267, z=-3.091, p=0.002). We see a non-significant decrease in match to reference for human subjects when comparing the Numeric condition to its compositional counterpart, with performance dropping from approximately 0.6 to approximately 0.5 (coef=-0.3795, z=-1.028, p=0.304).
By contrast, we do not find the models to be significantly worse in compositional conditions than non-compositional ones. In fact, GPT-4 exhibits a slight improvement in the compositional conditions, though this change is statistically insignificant for both the Dual-attribute condition relative to the Categorial condition (coef=0.2281, z=0.477, p=0.634) and for the Numeric Dual-attribute condition relative to the Numeric condition (coef= 0.1292, z=0.254, p=0.799). For Claude 3 we similarly find the differences to be insignificant for the Dual-attribute condition relative to the Categorial condition (coef=0.2049, z=0.452, p=0.651) and for the Numeric Dual-attribute condition relative to the Numeric condition (coef=-0.4013, z=-0.893, p=0.372). Llama-405B is insignificantly worse in the Dual-attribute condition relative to the Categorial condition (coef=-0.1030, z=-0.227, p=0.821), and its anomalously low performance in the Numeric Dual-attribute condition yields a significant difference in that condition relative to the Numeric condition (coef=-1.4323, z=-2.622, p=0.009).
@@ 4.2.4 Takeaways
The Semantic Content experiment confirms that human subjects perform robustly and flexibly across diverse task variations. Claude 3 matches human performance in all conditions, indicating it shares humans’ tendency to use the source domain’s semantic content when completing target domains. While GPT-4’s poor performance in numeric conditions is notable, it most likely reflects a failure in numeric reasoning rather than a difference in analogical reasoning.
We find evidence of decreased human performance, but not model performance, in compositional conditions, contrasting with some existing research [36]. However, other factors may be at play. Models’ negative compositionality effect may be masked by a positive effect, such as increased available information: when the target domain represents two source domain properties, models may more easily recognize the encoding of source domain properties. Human subjects may benefit less from this competing effect if they do not struggle to observe this information.
4.3 Evaluating the effect of chain-of-thought
We provide the best-performing model, Claude 3, with a chain-of-thought prompt across all experiments to investigate whether this improves performance. Across several prompt variations we observe weak results, documented in A.3. Therefore, the prompting strategy reported in the main text does not appear to be uncharitable to models.
@@ 5 Discussion
Our results show that the best-performing LLMs are able to successfully complete many analogical reasoning tasks in a way that is comparable to humans, even when using novel stimuli that are highly unlikely to appear in their training data. That said, behavior variation across task presentations and control conditions reveals that there remain meaningful differences in how such analogies are processed, evidenced by differences in how humans and models respond to distracting or misleading information. However, we observe a clear trend in which more recent leading models come increasingly close to matching human performance across our tasks. In particular, Claude 3 exhibits impressively robust performance across most task variations, even closing the gap with humans in some test conditions in which the earlier GPT-4 exhibited limitations (such as the $2\times n$ task version in which mapping from the source domain must be used for success). Within our sample of models, GPT-4 and Claude 3 may have distinctive features correlated with their recency, such as size or innovations in architecture and post-training, though such features are difficult to judge given that the models are closed-source. Within the open-source models, the good performance of the very large Llama 405B despite a relatively straightforward design (it is not a mixture of experts and undergoes standard post-training approaches) suggests that parameter and data scale may underlie a significant but unknown proportion of recent performance increases, although other factors are likely implicated. Together, these results raise questions about the ability of LLMs and similar models to serve as candidate cognitive models, which we discuss briefly below.
@@ 5.1 Evaluating the Competence of LLMs
The breadth of Claude 3’s success in our tasks - that likely do not appear in its training data - is noteworthy. It suggests that state-of-the-art LLMs can broadly match human performance not only in formal analogical reasoning tasks but also in tasks that require re-representing and transferring semantic information across linguistic and non-linguistic domains. As such, our results weigh against a view held by some cognitive scientists according to which connectionist models without a built-in symbolic component are constitutively limited in their ability to robustly handle analogical reasoning tasks. This view is bolstered by a long history of only modest progress in analogical reasoning using deep learning systems, coupled with criticisms of their limited successes as being overly fit to data that should be treated more appropriately as withheld generalization test sets [37]. They also inform discussions of whether LLMs possess “functional” linguistic competence, in addition to “formal” linguistic competence [38]. Further work is needed to characterize the precise mechanism that LLMs are using to solve these tasks; it is possible–though increasingly unlikely given the robustness of the behavioral results–that success is due to a myriad of heuristics rather than a systematic analogical reasoning process. Even so, evidence of LLMs completing analogical reasoning tasks in domains designed to involve relating semantic information to non-linguistic correlates, in addition to tasks over abstract symbols, supports the claim that some LLMs are capable of functional linguistic competence in addition to formal competence. Such an outcome would be consistent with the argument made in Mahowald et al. [38], which left room for the development of such capabilities from particular types of post-training to complement next-word prediction, and all successful models we identify undergo significant post-training.
@@
From this perspective, our results offer evidence to support both sides of the present debate about whether LLMs possess human-level analogical reasoning (see Webb et al. [10], Lewis and Mitchell [9], and Hodel and West [47]). Supporting the argument of Webb et al. [10] that deficiencies in capabilities other than analogical reasoning can explain poor model performance in some tasks, we find that GPT-4’s failure in the numeric conditions of our Semantic Content experiment may be due to a deficiency in counting ability. However, contrary to Webb et al. [7], who report impressive analogical reasoning in both GPT-3 and GPT-4, we do find a notable difference in the performance of these two models, with GPT-3 performing quite poorly on our tasks. Among the models tested, only GPT-4, Claude 3, and Llama-405B produce results that merit detailed comparison with human subjects. This suggests that claims of human-level performance of LLMs on analogical reasoning tasks may have been premature and might have relied on insufficiently challenging tasks.
@@
Collectively, these patterns bear on the larger question of how we should arbitrate disputes about competence in machine-human comparisons. On the one hand, it seems reasonable to assume that any system that can reliably achieve success at or above human level on experiments like ours–without relying on memorization and other confounds–should be considered competent at analogical reasoning. On the other hand, we should be open to the possibility that such competence may be implemented differently in LLMs and humans.
@@
5.2 Can LLMs Support Progress in Understanding Analogical Reasoning?
A central question for our study and similar studies is: what is the significance of LLMs’ analogical reasoning abilities to human cognitive theory? By designing a set of analogical reasoning tasks which are not readily explained by existing theories and models of analogy, we demonstrate that LLMs, at a minimum, expand the empirical coverage of computational models of analogy. This finding alone necessitates that they be taken seriously, and that their mechanisms be investigated further, before their significance to human analogical reasoning is dismissed.
More specifically, we can consider how our tasks relate to one of the leading theories of analogical reasoning, structure mapping theory (SMT). A task that is directly covered by SMT involves establishing the correspondence between two domains based on aligning higher-order relations that ignore the semantic content of the entities of each domain. By contrast, our experiment 1) involves semantic content in one of the two experiments, 2) requires subjects to make an analogical inference based on a given mapping by performing rule induction [17] or schema induction [48] to make inferences about the target domain, and 3) requires the flexible re-representation of object terms so as to foreground the analogy-relevant properties of the domains, such as when foregrounding the properties of being a mammal and having $n$ legs when solving one of our problems (whereas most of the many possible concept representations do not highlight salient features and therefore do not support inference). Each of these three aspects corresponds to an open set of theoretical problems in understanding analogical reasoning.
@@ 5.2 Can LLMs Support Progress in Understanding Analogical Reasoning?
Incorporating semantic information into our analogical reasoning problems ensures that we investigate whether LLMs’ analogical reasoning resembles that of human subjects in a relevant way to its purportedly central role in broader cognition. Following Gentner [13], emphasis has been placed on relational similarity, rather than just feature similarity, in mapping from a familiar source to a foreign target domain during analogical reasoning to allow for the flexible transfer of knowledge [49, 50, 51]. This conception allows analogical reasoning to play a fundamental role in human cognition, supporting the emergence of diverse cognitive abilities via “bootstrapping” [52, 53, 54]. In bootstrapping, two cognitive processes mutually support each other’s development. In Gentner’s Structure-Mapping Theory (SMT), language development and structure-mapping-based analogical reasoning are hypothesized to co-develop, with structure-mapping developing the necessary relational reasoning to model language-world relations, and language acquisition in turn developing symbolic reasoning capacities that amplify structure-mapping abilities. Consequently, analogical reasoning is seen as a central cognitive phenomenon of interest.
Implicitly requiring subjects in our experiments to infer an abstract rule or schema that governs the relationship between a source and target domain, and to do so based on flexibly re-representing inputs in a suitable way to support this rule discovery, addresses the core of an open problem in analogical reasoning. As Chalmers et al. argue, analogical reasoning over inputs that are already symbolically represented in a manner that supports such reasoning can be trivial and not informative of analogical inputs over unconstrained natural inputs: problem-specific representations hand-crafted by researchers can offload a significant portion of the reasoning problem [11]. Before a subject can search over a large space of possible rules governing the correspondence between two domains, they must first search over a large space of possible ways of representing the entities in the domains, only some of which correspond to simple rules relating the domains. This feature-selection process requires non-trivial reasoning about which properties of the system are relevant in the given context.
The requirement of our tasks to perform rule or schema induction based on the re-representation of real-world inputs is important for understanding analogical reasoning’s role in natural contexts. Current theories either address re-representation without real-world concepts, or handle real-world concepts without tackling re-representation. For example, Mitchell and Hofstadter’s Copycat models [14] addressed re-representation but in the domain of letter string analogies, while subsequent work addressed re-representation in a synthetic Raven’s matrices domain [16]. Conversely, recent efforts by Keith and Hongjing [15] [55] address real-world inputs but neglect the problem of re-representation.
The success of some LLMs in many of our tasks suggests that the most advanced models may be capable of employing an analogical reasoning approach that captures some aspects of human analogical reasoning in a domain that is not well-covered by existing theory. The potential emergence of such competence from training primarily on text prediction could yield new hypotheses about the emergence of analogical reasoning as a central cognitive faculty from generic learning mechanisms (possibly combined with the unique pressures of language acquisition). [^5] While we do not know the proportion of simple text prediction in the training of closed models such as GPT-4 and Claude 3, the bulk of the relatively-successful Llama 405B’s training appears to be text prediction. The model is pre-trained on predicting 15.6 trillion text tokens, while the scale of post-training appears significantly smaller - see sections 4.2 and 4.3 of [26]. That said, the majority of Llama 405B’s training being text prediction does not guarantee that this is what underlies the development of its analogical reasoning capabilities. The proprietary nature of leading LLMs like Claude 3 limits our ability to directly investigate the features that may explain the emergence of a response pattern largely mirroring that of human subjects. However, the open-weights Llama-405B comes tantalizingly close to replicating the performance of Claude 3 on our tasks. This represents a significant advance for the field of cognitive science, as subsequent development may provide an open computational model capable of reproducing non-trivial human analogical reasoning. The study of analogical reasoning may therefore be approaching an era in which available theories are partial but in which a good empirical computational model exists, allowing for future interpretability work on LLMs to support the refinement of cognitive theories of analogy.
@@
That said, the mixed success of LLMs and the significant differences from humans in certain conditions underscore the need for continued research to test the robustness of any conclusion that analogical reasoning in LLMs closely matches that of human subjects. As LLM outputs continue to converge toward human responses–an expected product of the language modeling objective–it is crucial to develop novel tasks that examine analogical reasoning ability and are almost certainly not attested in the training data. While our task allows for clear discrimination between human performance and that of most models prior to Claude 3, further differences in analogical reasoning patterns between humans and Claude 3 likely exist beyond those revealed by our tests. More granular testing would help clarify the extent of the remaining discrepancies between humans and the most advanced LLMs.